Assume that females have pulse rates that are normally distributed with a mean mu= 72.0 beats per minute and a standard deviation of alpha =12.5 beats per minute. Complete parts​ (a)
a. If 1 adult female is randomly​ selected, find the probability that her pulse rate is less than 79 beats per minute. Round to four decimal place
1
Expert's answer
2020-11-09T20:08:19-0500

"P(x<79)=P(z<\frac{79-72}{12.5})\\
=P(z<0.56)=0.5+P(0<z<0.56)\\
=0.5+0.2123=0.7123"